Can you please help me with this problem?
Myrna is 27 years older than her son.
A year ago, she was twice as old as him.
how old is Myrna now?

Answer by stanbon(75887)

Myrna is 27 years older than her son.
A year ago, she was twice as old as him.
how old is Myrna now?
Equations:
M = s + 27
M-1 = 2(s-1)
------
Substitute for "M" and solve for "s":
s+26-1 = 2(s-1)
s+25 = 2s-2
s = 27 (son's age now)
--------
Since Myrna = s + 27, Myrna = 27+27 = 54 (Myrna's age now)
===============================================================
Cheers,
Stan H.